OpenStudy (anonymous):

Write the equation of the line that passes through (-4, 12) and (7, 23). A. y = x + 16 B. y = 3x + 18 C. y = 1/6 x + 16 D. y = 1/3 x + 18

OpenStudy (anonymous):

23-12=11 7- -4= 11 11 divided by 11 is 1 (slope) 12=1(-4)+b +4 +4 16=b y=x+16

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no prob!